OpenStudy (anonymous):

In 18.01 Excercises by Prof. David Jerison unit 1 Differentiation 1A-3 how do you identify a function as odd,even or periodic

OpenStudy (anonymous):

A)\[x ^{3}+3x/1-x ^{4}\] B)\[\sin^{2}x \] C)\[\tan x/1+x ^{2}\] D)\[(1+x)^{4}\]

OpenStudy (jkristia):

http://en.wikipedia.org/wiki/Even_and_odd_functions if f(x) = f(-x) then the graph has symmetry with y axis and is an even function (the power of the function is even, e.g. x^2) if -f(x) = f(-x) then the graph has symmetry around origin and the function is and odd (the power of the function will alsy be odd, e.g. x^3). Not sure about the periodic other than the function contains one or more of the periodic functions.

OpenStudy (anonymous):

thanks,i see that if \[f(x+p)=f(x)\] for all x values in a set of domain,where p is a positive constant,then f is called a Periodic Function and the smallest such number p is called a period.For instance \[y=sinx\] has a period \[2pi\] so trigonometric graphs are periodic graphs. Thanks again...